OpenStudy (anonymous):

2nd integer question

OpenStudy (anonymous):

Go ahead

OpenStudy (anonymous):

It says that the integers are consecutive, so each integer must be 1 more than the integer before it. So, the table will look exactly like the one in the previous question. OK?

OpenStudy (anonymous):

In this question, you need to add the 2nd and 3rd integers. What do you get?

OpenStudy (anonymous):

x+1 then x+3

OpenStudy (anonymous):

Sorry, I messed up. It says that the integers are consecutive EVEN integers. So if the first EVEN integer is x, what is the next EVEN integer?

OpenStudy (anonymous):

x+1?

OpenStudy (anonymous):

Not quite. For example, if the first even integer is 4, what is the next even integer?

OpenStudy (anonymous):

x+5

OpenStudy (anonymous):

Whoa. Let's back up. What is the next largest even integer after 4?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

Do you know what an even number is?

OpenStudy (anonymous):

It's OK. Yes or no?

OpenStudy (anonymous):

The next even number after 4 is 6

OpenStudy (anonymous):

Correct. So, if you have any even integer, how much to you have to add to it to get the next largest even integer?

OpenStudy (anonymous):

the next even number from x would be 2 correct?

OpenStudy (anonymous):

No. If x is an even integer, how much do you have to add to it to get the next largest even integer?

OpenStudy (anonymous):

It seems you recognize that even integers are two numbers apart. So the next largest even integer to an even integer x is 2 more than x. How do you write "2 more than x" mathematically?

OpenStudy (anonymous):

Hello? Do you still help?

OpenStudy (anonymous):

*need help?

OpenStudy (anonymous):

I'm sorry I'm just really stressed out on this assignment, I've been doing it for over 3 hours.

OpenStudy (anonymous):

OK. Stay with me and we'll work it out together. Can you answer my last question?

OpenStudy (anonymous):

2+x

OpenStudy (anonymous):

Right! Except we would normally write it as x + 2. What is the next even integer after x + 2 ?

OpenStudy (anonymous):

x+4

OpenStudy (anonymous):

Excellent. Now your table is filled out. First even integer: x Next even integer: x + 2 Next even integer: x + 4 The question asks for the sum of the second and third ones. Can you do that?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

I think I did it wrong. I got 6.

OpenStudy (anonymous):

\[\left( x+2 \right) + \left( x+4 \right) = ?\]

OpenStudy (anonymous):

2x+6

OpenStudy (anonymous):

Correct. Good job!